# Green currents of holomorphic correspondences on compact Kähler manifolds

### 06 Mar 2026 v1

@@ 5. Examples
In this section, we prove that the class of holomorphic correspondences with small inverse multiplicity is very rich and contains many interesting examples.
@@ 5.1. Families of holomorphic correspondences
When $k=d$, the result follows directly from Lemma 3.2: since the proof uses a local argument and we only need to consider what happens in a neighbourhood of $y$, we do not need $Y$ to be compact.
@@
Let $\mathbb{D}_{k-d}$ and $\mathbb{D}_{Y}$ be the poly-disks centered at 0 in, respectively, $\mathbb{C}^{k-d}$ and $Y$ such that $\mathbb{D}_{J}=\mathbb{D}_{k-d}\times\mathbb{D}_{Y}$. For $(x_{J},y)\in\mathbb{D}_{J}$, denote by $m_{(x_{J},y)}$ the measure defined by the sum of the Dirac measures given by the points in the fibre $\pi_{J}|_{{\Gamma}\cap\mathbb{D}_{X\times Y}}^{-1}(x_{J},y)$, counted with multiplicity. By Lemma 3.2, for fixed $x_{J}$, when $y_{i}$ converges to $y$ in $\mathbb{D}_{Y}$, we have $m_{(x_{J},y_{i})}\to m_{(x_{J},y)}$. Now for any $y\in\mathbb{D}_{Y}$, the restricted projection map $\pi|_{\Gamma_{y}\cap\mathbb{D}_{X\times Y}}:\Gamma_{y}\cap\mathbb{D}_{X\times Y}\to\mathbb{D}_{k-d}$ is still proper. Therefore, it is also a ramified covering. For $x_{J}\in\mathbb{D}_{k-d}$ generic, $m_{(x_{J},y_{i})}$ equals to the slicing of $\Gamma_{y}]$ with respect to $\pi|_{\Gamma_{y}\cap\mathbb{D}_{X\times Y}}$ at the point $x_{J}$. From the case $k=d$ and the continuity of the slice with respect to the parameter [36, Theorem 3.3.2], the same is true for every $x_{J}\in\mathbb{D}_{k-d}$. Let $\varphi$ be a smooth test function compactly supported in $\mathbb{D}_{X\times Y}$. Denote by $\omega^{J}$ the volume form on $\mathbb{D}_{k-d}$. Then, by the bounded convergence theorem, we have
$\lim_{i\to\infty}\langle[\Gamma_{y_{i}}]\wedge\omega^{J},\varphi\rangle$ $=\lim_{i\to\infty}\int_{\mathbb{D}_{k-d}}\langle m_{(x_{J},y_{i})},\varphi\rangle\omega^{J}(x_{J})$
$=\int_{\mathbb{D}_{k-d}}\langle m_{(x_{J},y)},\varphi\rangle\omega^{J}(x_{J})=\langle[\Gamma_{y}]\wedge\omega^{J},\varphi\rangle.$
@@
    there is an analytic subset $\Gamma$ of pure dimension $k+k^{\prime}$ on $\Sigma\times X\times X$ such that $\langle[\Gamma],\pi_{\Sigma},t\rangle=[\Gamma_{t}]$ for every $t$. It is called the graph of the family.
@@
We are going to prove that having small inverse multiplicity is a generic property for holomorphic correspondences. The next result implies that for a family of holomorphic correspondences $\{f_{t}\}_{t\in\Sigma}$ on $X$, the condition about the local multiplicity $\rho(f_{t})<\rho$ is true for generic $t$ as long as the set $\Sigma_{\rho}:=\{t\in\Sigma:\rho(f_{t})<\rho\}$ is non-empty. When also $\{f_{t}^{-1}\}_{t\in\Sigma}$ is a family of holomorphic correspondences, since $\delta(f_{t})=\rho(f_{t}^{-1})$, the same result holds for the inverse multiplicity.
@@
Notice that $\pi_{\Sigma}(W)=\Sigma\setminus\Sigma_{\rho}$. Therefore, it suffices to prove that $W$ is an analytic set. Notice that this is a local property. For any point $(t_{0},x_{0},y_{0})\in\Gamma$, by the proof of Lemma 3.2, we can choose two neighbourhoods $U_{0}$ of $(t_{0},x_{0})$ and $V_{0}$ of $y_{0}$ such that $\pi_{1}|_{\Gamma\cap(U_{0}\times V_{0})}$ is a ramified covering of degree $d$. Here $d$ is smaller than or equal to $d_{0}(f_{t})$. We may choose $U_{0}$ and $V_{0}$ as local coordinate charts with local coordinates $(t^{1},\ldots,t^{k^{\prime}};x^{1},\ldots,x^{k})$ and $(y^{1},\dots,y^{k})$. For any $a\in U_{0}$, denote by $a_{1},\dots,a_{d}$ the preimages of $a$ under $\pi_{1}|_{\Gamma\cap(U_{0}\times V_{0})}$, counted with multiplicity. Define functions
@@ 5.2. Symmetric product and polynomial correspondences
Lemma 5.6. Let $h$ be a holomorphic correspondence on a compact Riemann surface $X$ whose graph contains no fiber of $\pi_{1}$ or $\pi_{2}$. Suppose the critical values of $h$ are not periodic. Then there exists an integer $D>0$ such that $\delta(h^{n})\leq D$ for all $n$.
@@
    If $\widehat{f}^{-1}$ is a holomorphic correspondence, then so is $f^{-1}$. In this case, the inverse multiplicity of $f^{N}$ satisfies $\delta(f^{N})\leq k!\cdot\delta(\widehat{f}^{N})$ for all $N$;
@@
    If $h$ is given by $\Gamma_{h}=\{(z,w)\in\mathbb{P}^{1}\times\mathbb{P}^{1}:g_{1}(z)=g_{0}(w)\}$ where $g_{i}$ is a holomorphic endomorphism of $\mathbb{P}^{1}$ of degree $d_{i}$ for $i=0$ and $1$, then $f$ is given by $\Gamma_{f}=\{(Z,W)\in\mathbb{P}^{k}\times\mathbb{P}^{k}:G_{0}(Z)=G_{1}(W)\}$ where $G_{i}$ is a holomorphic endomorphism of $\mathbb{P}^{k}$ of degree $d_{i}$ for $i=0$ and $1$.
@@
$G_{0}(\pi(z_{1},\dots,z_{k}))=\pi(g_{0}(z_{1}),\dots,g_{0}(z_{k}))=\pi(g_{1}(w_{1}),\dots,g_{1}(w_{k}))=G_{1}(\pi(w_{1},\dots,w_{k}))$
@@
Corollary 5.11. For $0<q\leq k$ and $d_{1}>d_{0}$, there exists an integer $N$ such that for generic elements $f$ in $\mathcal{F}_{(d_{0},d_{1})}$, $f^{N}$ has $q$-small inverse multiplicity.
@@
for any $n\geq 1$. Notice that for any $f\in\mathcal{F}_{(d_{0},d_{1})}$ and $n\geq 1$, $d_{q}(f^{n})/d_{q-1}(f^{n})=(d_{1}/d_{0})^{n}$. Choose an integer $N$ such that $25k(4k!D^{k})^{k-q+1}<(d_{1}/d_{0})^{N}$. Apply Proposition 5.1 to the family $\{f^{-N}\}_{f\in\mathcal{F}_{(d_{0},d_{1})}}$ and $\rho=\frac{1}{4}[\frac{1}{25k}(\frac{d_{1}}{d_{0}})^{N}]^{1/(k-q+1)}$. Notice that $\Sigma_{\rho}$ is exactly the set of $f\in\mathcal{F}_{(d_{0},d_{1})}$ such that $f^{N}$ has $q$-small inverse multiplicity. Then we have $f_{0}\in\Sigma_{\rho}$, and we deduce that $\Sigma_{\rho}$ is a Zariski dense open set, which is the desired result. ∎

### 25 Mar 2026 v2

@@ 5. Examples
In this section, we prove that the class of holomorphic correspondences with small adjoint multiplicity is very rich and contains many interesting examples.
@@ 5.1. Families of holomorphic correspondences
When $k=d$, the result follows directly from Lemma 3.2 applied with $X_{1}=Y$, $X_{2}=X$, and $A=\Gamma$. Since the proof of the Lemma uses a local argument and we only need to consider what happens in a neighbourhood of $y$, we do not need $Y$ to be compact.
@@
Let $\mathbb{D}_{k-d}$ and $\mathbb{D}_{Y}$ be the poly-disks centered at 0 in, respectively, $\mathbb{C}^{k-d}$ and $Y$ such that $\mathbb{D}_{J}=\mathbb{D}_{k-d}\times\mathbb{D}_{Y}$. For $(x_{J},y)\in\mathbb{D}_{J}$, denote by $m_{(x_{J},y)}$ the measure defined by the sum of the Dirac measures given by the points in the fibre $\pi_{J}|_{{\Gamma}\cap\mathbb{D}_{X\times Y}}^{-1}(x_{J},y)$, counted with multiplicity. By Lemma 3.2, for fixed $x_{J}$, when $y_{i}$ converges to $y$ in $\mathbb{D}_{Y}$, we have $m_{(x_{J},y_{i})}\to m_{(x_{J},y)}$. Now for any $y\in\mathbb{D}_{Y}$, the restricted projection map $\pi|_{\Gamma_{y}\cap\mathbb{D}_{X\times Y}}:\Gamma_{y}\cap\mathbb{D}_{X\times Y}\to\mathbb{D}_{k-d}$ is still proper. Therefore, it is also a ramified covering. For $x_{J}\in\mathbb{D}_{k-d}$ generic, $m_{(x_{J},y_{i})}$ equals to the slicing of $\Gamma_{y}]$ with respect to $\pi|_{\Gamma_{y}\cap\mathbb{D}_{X\times Y}}$ at the point $x_{J}$. From the case $k=d$ and the continuity of the slice with respect to the parameter [36, Theorem 3.3.2], the same is true for every $x_{J}\in\mathbb{D}_{k-d}$. Let $\varphi$ be a smooth test function compactly supported in $\mathbb{D}_{X\times Y}$. Denote by $\omega_{J}$ the volume form on $\mathbb{D}_{k-d}$. Then, by the bounded convergence theorem, we have
$\lim_{i\to\infty}\langle[\Gamma_{y_{i}}]\wedge\omega_{J},\varphi\rangle$ $=\lim_{i\to\infty}\int_{\mathbb{D}_{k-d}}\langle m_{(x_{J},y_{i})},\varphi\rangle\omega_{J}(x_{J})$
$=\int_{\mathbb{D}_{k-d}}\langle m_{(x_{J},y)},\varphi\rangle\omega_{J}(x_{J})=\langle[\Gamma_{y}]\wedge\omega_{J},\varphi\rangle.$
@@
    there is an effective analytic cycle $\Gamma$ of pure dimension $k+k^{\prime}$ on $\Sigma\times X\times X$ such that $\langle[\Gamma],\pi_{\Sigma},t\rangle=[\Gamma_{t}]$ for every $t$. It is called the graph of the family.
@@
We are going to prove that having small adjoint multiplicity is a generic property for holomorphic correspondences. The next result implies that for a family of holomorphic correspondences $\{f_{t}\}_{t\in\Sigma}$ on $X$, the condition about the local multiplicity $\rho(f_{t})<\rho$ is true for generic $t$ as long as the set $\Sigma_{\rho}:=\{t\in\Sigma:\rho(f_{t})<\rho\}$ is non-empty. When also $\{f_{t}^{-1}\}_{t\in\Sigma}$ is a family of holomorphic correspondences, since $\delta(f_{t})=\rho(f_{t}^{-1})$, the same result holds for the adjoint multiplicity.
@@
Notice that $\pi_{\Sigma}(W)=\Sigma\setminus\Sigma_{\rho}$. Therefore, it suffices to prove that $W$ is an analytic set. Notice that this is a local property. Since $\pi_{1}|_{\Gamma}$ has finite fibres, it is a ramified covering. So, for every point $(t_{0},x_{0},y_{0})\in\Gamma$ we can choose two neighbourhoods $U_{0}$ of $(t_{0},x_{0})$ and $V_{0}$ of $y_{0}$ such that $\pi_{1}|_{\Gamma\cap(U_{0}\times V_{0})}$ is a ramified covering of degree $d$. Here $d$ is smaller than or equal to $d_{0}(f_{t})$. We may choose $U_{0}$ and $V_{0}$ as local coordinate charts with local coordinates $(t^{1},\ldots,t^{k^{\prime}};x^{1},\ldots,x^{k})$ and $(y^{1},\dots,y^{k})$. For every $a\in U_{0}$, denote by $a_{1},\dots,a_{d}$ the preimages of $a$ under $\pi_{1}|_{\Gamma\cap(U_{0}\times V_{0})}$, counted with multiplicity. Define functions
@@ 5.2. Symmetric product and polynomial correspondences
Lemma 5.6. Let $h$ be a holomorphic correspondence on a compact Riemann surface $X$ whose graph contains no fibre of $\pi_{1}$ or $\pi_{2}$. Suppose the critical values of $h$ are not periodic. Then there exists an integer $D>0$ such that $\delta(h^{n})\leq D$ for all $n$.
@@
    If $\widehat{f}^{-1}$ is a holomorphic correspondence, then so is $f^{-1}$. In this case, the adjoint multiplicity of $f^{N}$ satisfies $\delta(f^{N})\leq k!\cdot\delta(\widehat{f}^{N})$ for all $N$;
@@
    If $h$ is given by $\Gamma_{h}=\{(z,w)\in\mathbb{P}^{1}\times\mathbb{P}^{1}:g_{0}(z)=g_{1}(w)\}$ where $g_{i}$ is a holomorphic endomorphism of $\mathbb{P}^{1}$ of algebraic degree $d_{i}$ for $i=0$ and $1$, then $f$ is given by $\Gamma_{f}=\{(Z,W)\in\mathbb{P}^{k}\times\mathbb{P}^{k}:G_{0}(Z)=G_{1}(W)\}$ where $G_{i}$ is a holomorphic endomorphism of $\mathbb{P}^{k}$ of algebraic degree $d_{i}$ for $i=0$ and $1$.
@@
$G_{0}(\pi(z_{1},\dots,z_{k}))$ $=\pi(g_{0}(z_{1}),\dots,g_{0}(z_{k}))$
$=\pi(g_{1}(w_{1}),\dots,g_{1}(w_{k}))=G_{1}(\pi(w_{1},\dots,w_{k}))$
@@
Corollary 5.11. For $0<q\leq k$ and $d_{1}>d_{0}$, there exists an integer $N$ such that for generic elements $f$ in $\mathcal{F}_{(d_{0},d_{1})}$, $f^{N}$ has $q$-small adjoint multiplicity.
@@
for any $n\geq 1$. Notice that for any $f\in\mathcal{F}_{(d_{0},d_{1})}$ and $n\geq 1$, $d_{q}(f^{n})/d_{q-1}(f^{n})=(d_{1}/d_{0})^{n}$. Choose an integer $N$ such that $25k(4k!D^{k})^{k-q+1}<(d_{1}/d_{0})^{N}$. Apply Proposition 5.1 to the family $\{f^{-N}\}_{f\in\mathcal{F}_{(d_{0},d_{1})}}$ and $\rho=\frac{1}{4}[\frac{1}{25k}(\frac{d_{1}}{d_{0}})^{N}]^{1/(k-q+1)}$. Notice that $\Sigma_{\rho}$ is exactly the set of $f\in\mathcal{F}_{(d_{0},d_{1})}$ such that $f^{N}$ has $q$-small adjoint multiplicity. Then we have $f_{0}\in\Sigma_{\rho}$, and we deduce that $\Sigma_{\rho}$ is a Zariski dense open set, which is the desired result. ∎